Four particles having masses m, 2m, 3m and 4m are placed at the four corners of a square of edge a. Find the gravitational force acting on a particle of mass m placed at the centre.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. GM/(R + h)^2 = 6.67 * 10^-11 * 7.4 * 10^22/(1740 + 1000)^2 = 49.358 * 10^11/2740 * 2740 * 10^6 = 49.358 * 10^11/0.75 * 10^13 = 65.8 * 10^-2 = 0.65 m/s^2